OpenStudy (anonymous):

SOMEONE PLEASE HELP THIS NEEDS TO BE FINISHED TODAY! A(t)=P(1+r/n)^nt A(t)=39,145(1+.03/12)^12t Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal. Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate.

OpenStudy (anonymous):

(Being a smart) Is question 1, question 2 is (Lastly, you decide)

OpenStudy (the_fizicx99):

Did you write that function?

OpenStudy (anonymous):

No the function came like that in the assignment.

OpenStudy (anonymous):

Heres the practice problem from the lesson that was most similar to the assignment.

OpenStudy (the_fizicx99):

This is Algebra II ._. But from my understanding, 10,000 = 39,145(1 +\(\ \dfrac{0.03}{n}\))\(^{nt}\) I'm not familiar with Logarithm yet. So I don't know how to solve for \(\ t \) yet :/ maybe @ganeshie8 can help you

OpenStudy (anonymous):

@tJe_FiZiCx99 Thanks for the assistance though, it is algebra II with a logarithm. Something about using the Power Property of logarithms didn't exactly help with the bases being unequal.

OpenStudy (anonymous):

@tHe_FiZiCx99 *

OpenStudy (the_fizicx99):

Sorry, I'm taking Alg II and Geometry in a few weeks though. See you're in flvs though, have fun!

ganeshie8 (ganeshie8):

yes, just set it equal to 10,000 and solve t

OpenStudy (anonymous):

@tHe_FiZiCx99 Haha thanks, I'll try to have "fun" with it xD. @ganeshie8 I was having trouble with that.

OpenStudy (the_fizicx99):

cx

ganeshie8 (ganeshie8):

\(\large 10,000 =39,145(1+\frac{.03}{12})^{12t} \)

ganeshie8 (ganeshie8):

start by dividing 39,145 both sides : \(\large \frac{10,000}{39,145} = (1+\frac{.03}{12})^{12t} \)

ganeshie8 (ganeshie8):

\(\large 0.25546= (1+\frac{.03}{12})^{12t} \)

ganeshie8 (ganeshie8):

now take log both sides

OpenStudy (anonymous):

@ganeshie8 I got stuck after dividing on the first time I attempted this, I don't really understand by what you mean taking log on both sides though.

ganeshie8 (ganeshie8):

\(\large 0.25546= (1+\frac{.03}{12})^{12t} \) \(\large 0.25546= (1.0025)^{12t} \) take log both sides \(\large \log 0.25546= \log (1.0025)^{12t} \)

ganeshie8 (ganeshie8):

next use below property : \(\large \log a^b = b \log a\)

ganeshie8 (ganeshie8):

\(\large 0.25546= (1+\frac{.03}{12})^{12t} \) \(\large 0.25546= (1.0025)^{12t} \) take log both sides \(\large \log 0.25546= \log (1.0025)^{12t} \) \(\large \log 0.25546= 12t \log (1.0025) \) \(\large \frac{\log 0.25546}{12\log (1.0025) }= t \)

ganeshie8 (ganeshie8):

use ur calculator to evaluate the left hand side

OpenStudy (anonymous):

@ganeshie8 Thanks a bunch! But what about the last part, #3?